# Transformers as Implicit State Estimators: In-Context Learning in Dynamical Systems

### v2

@@ 1 Introduction
In-context learning (ICL), particularly in the form of few-shot prompting [26], has emerged as a prominent research direction in natural language processing (NLP). In this framework, a large language model (LLM) performs new tasks by conditioning on a small number of input–output examples provided in the prompt. One of the earliest works to demonstrate this capability was by Brown et al. [3], who evaluated GPT-3 across a range of NLP datasets under “zero-shot", “one-shot" and “few-shot" settings. Zhao et al. [28] identified majority-label bias, recency bias, and common-token bias as sources of instability in GPT-3’s accuracy under few-shot prompting, and proposed contextual calibration as a remedy. A theoretical explanation for ICL as a form of implicit Bayesian inference is offered in Xie et al. [25]. Min et al. [14] explore why ICL works in practice, showing that its success is not dependent on having access to ground-truth labels; it further emphasized the importance of label space, input distribution, and prompt format. In parallel, Schlag et al. [20] provide a theoretical analysis showing that transformers can act as fast weight programmers.
@@
Early work on using standard transformer decoders for in-context learning of autoregressive models includes [5], where the ability of transformers to learn function classes from examples is empirically investigated. A model is said to learn a function class $\mathcal{F}$ over a domain $\mathcal{X}$ if, for any $f\in\mathcal{F}$ and any inputs $x_{1},x_{2},\ldots,x_{N},x_{\text{query}}$ drawn i.i.d. from $\mathcal{X}$, the model can predict $f(x_{\text{query}})$ given the sequence $x_{1},f(x_{1}),x_{2},f(x_{2}),\ldots,x_{N},f(x_{N}),x_{\text{query}}$. The function classes studied by Garg et al. [5] range from simple linear functions to sparse linear models, two-layer neural networks, and decision trees.
@@
Two parallel works by Von Oswald et al. [23] and Akyürek et al. [1] examine which algorithms transformers implicitly implement when learning function classes in-context. Von Oswald et al. [23] build on Schlag et al. [20] and show that linear self-attention layers can be interpreted as performing a step of gradient descent. Specifically, they prove that for a single-headed linear attention layer, there exist key, query, and value matrices such that a forward pass corresponds to one step of gradient descent with an $\ell_{2}$ loss applied to each token. In contrast, Akyürek et al. [1] introduce a raw operator framework that supports operations such as matrix multiplication, scalar division, and memory-like read/write. They demonstrate that a single transformer head equipped with suitable key, query, and value matrices can approximate this operator, implying that transformers are, in principle, capable of performing linear regression via either stochastic gradient descent or in closed-form.
@@
Our work investigates whether transformers, using in-context learning (ICL), can learn to perform filtering in dynamical systems described by noisy state-space models. Specifically, we study whether a transformer, conditioned on a short context of past input–output pairs (and, optionally, system parameters), can predict the current output of the system. The model is pre-trained on synthetic trajectories generated by randomly sampled system parameters and evaluated without gradient-based updates at test time. Our goal is to understand which classical filtering algorithms the transformer’s behavior most closely resembles under this setup. To that end, we begin by examining the structure of the Kalman filter [11], the optimal linear estimator for systems with Gaussian noise, and ask whether its operations can be replicated by a transformer architecture. We show that Kalman filtering can indeed be expressed using operations that are readily implemented by a transformer, and provide both a proof-by-construction and empirical evidence that, when sufficiently scaled and provided with enough context, the transformer learns to emulate Kalman-like behavior. Notably, its performance remains strong even when some parameters, such as the state-transition matrix, are omitted from the context, suggesting robustness and an ability to infer missing information implicitly. We then turn our attention to nonlinear systems and empirically demonstrate that transformers can in-context learn to achieve output prediction accuracy comparable to that of classical nonlinear filters such as the Extended Kalman Filter and particle filters. While our analysis and experiments primarily focus on output prediction, we also include preliminary evidence that transformers can, in some cases, explicitly recover latent states, thus highlighting their potential for more general inference tasks.
@@ 1.1 Related work
Prior research at the interface of deep learning and Kalman filtering include Deep Kalman Filters [13] and KalmanNet [17], which learns the Kalman gain in a data-driven manner using a recurrent neural network. A follow-up study [18] extends this approach using gated recurrent units to estimate both the Kalman gain and noise statistics while assuming fixed system parameters. In contrast, our model is trained on data generated from randomly sampled system parameters, encouraging the transformer to learn the filtering procedure itself rather than memorize input–output mappings for a specific system. Other works have explored connections between attention mechanisms and structured state-space models (SSMs). Dao & Gu [4] reformulate SSM computations as matrix multiplications on structured matrices, drawing parallels with efficient attention variants. Sieber et al. [21] propose a unified dynamical systems framework encompassing attention, SSMs, RNNs, and LSTMs. However, neither of these works address state estimation, filtering, or in-context learning. Goel & Bartlett [6] show that softmax self-attention can approximate the Nadaraya–Watson kernel smoother, which in turn resembles the Kalman filter. In contrast, our work directly targets in-context learning for dynamical systems and builds on the raw operator framework of Akyürek et al. [1] to show that transformers can implement the exact operations needed for Kalman filtering, as supported by both theoretical constructions and empirical validation.
@@ 2.1 Transformers
Transformers, introduced by Vaswani et al. [22], are a class of neural network architectures that rely on an attention mechanism to capture relationships among elements in an input sequence. This mechanism enables transformers to model long-range dependencies and has been central to their success in sequence-to-sequence tasks. The experiments in this paper utilize the GPT-2 architecture [16], a decoder-only variant of the transformer.
@@
where $\lambda$ denotes layer normalization and $\sigma$ is the activation function. In our experiments, we use the Gaussian Error Linear Unit (GeLU) activation [8].
@@ 2.2 State-space models for linear dynamical systems
State-space models are widely used across machine learning [7], computational neuroscience [2], control theory [9], signal processing [10], and economics [27]. A central problem in many of these domains is the estimation of the latent state sequence $x_{t}$} and its functions (e.g., the system’s output) given noisy observations $y_{t}$} and the parameters of the model.
@@ 2.3 In-context learning in the absence of dynamics
  • Stochastic Gradient Descent. Initialize with $\hat{x}_{0}=\mathbf{0_{n\times 1}}$, and update recursively as
    $\hat{x}_{t}=\hat{x}_{t-1}-2\alpha(h_{t-1}\hat{x}_{t-1}^{T}h_{t-1}-h_{t-1}y_{t-1}),$ (8)
@@
  • Ordinary Least Squares (OLS). Form the matrix $\mathbf{\bar{H}}\in\mathbb{R}^{N\times n}$ with rows $h_{1},h_{2},\dots,h_{N}$, and let $\mathbf{\bar{Y}}=[y_{1},y_{2},...,y_{N}]^{T}$. The OLS estimate is
    $\hat{x}_{OLS}=(\mathbf{\bar{H}}^{T}\mathbf{\bar{H}})^{-1}\mathbf{\bar{H}}^{T}\mathbf{\bar{Y}}.$ (9)
@@
    $\hat{x}_{Ridge}=(\mathbf{\bar{H}}^{T}\mathbf{\bar{H}}+\lambda I_{n\times n})^{-1}\mathbf{\bar{H}}^{T}\mathbf{\bar{Y}},$ (10)
@@
A pioneering study that explored the ability of language models to learn linear functions and implement simple algorithms was presented by Garg et al. [5]. Building on that work, Akyürek et al. [1] examined whether GPT-2–based transformers can in-context learn the setting where $x_{t}=x$ as in (7). Specifically, they investigated which algorithms the transformer implicitly learns to implement when tasked with predicting $y_{N}$, given an input formatted as the matrix
@@
In Akyürek et al. [1], the transformer was trained on batches of examples constructed from randomly sampled latent states and measurement parameters. The authors observed that the model’s behavior depends on both the architecture size and the context length. For small models trained with short contexts, the transformer approximates stochastic gradient descent (SGD). For moderate context lengths (up to the dimensionality of the latent state) and moderately sized models, its behavior resembles ridge regression. Finally, with large architectures and context lengths exceeding the latent dimension, the transformer’s performance approaches that of ordinary least squares (OLS).
A major contribution of Akyürek et al. [1] was to theoretically demonstrate that transformers can approximate the operations required to implement SGD and closed-form regression. This was accomplished by introducing the RAW (Read–Arithmetic–Write) operator (see Appendix G), parameterized by matrices $W_{o}$, $W_{a}$, and $W$, along with an element-wise operator $\circ\in[+,*]$. The RAW operator maps the input to layer $l$, denoted $\mathbf{q}^{l}$, to the output $\mathbf{q}^{l+1}$ using index sets $s$, $r$, $w$, a time set map $K$, and token positions $i=1,\dots,2N$ according to
$q_{i,w}^{l+1}$ $=W_{o}\left(\ \left[\frac{W_{a}}{|K(i)|}\sum_{k\in K(i)}q_{k}^{l}[r]\right]\circ Wq_{i}^{l}[s]\ \right),$ (11)
$q_{i,j\notin w}^{l+1}$ $=q_{i,j\notin w}^{l}.$ (12)
A single transformer head can approximate this operator for arbitrary choices $W_{o}$, $_{a}, WW$, and $\circ$. Moreover, there exist specific parametrizations $W_{o},W_{a},W,\circ\in\{+,*\}$ that allow the transformer to emulate key operations needed for SGD and closed-form regression, including affine transformations, matrix multiplication, scalar division, dot products, and memory-based read/write operations. For completeness, a detailed exposition of the RAW operator is provided in AppendixG.
@@ 3 In-Context Learning for Dynamical Systems
We begin by outlining an in-context learning procedure for the generic linear state-space model defined in (4)–(5), assuming a time-invariant state equation (i.e., $F_{t}=F\neq I$, $Q\neq 0$). For simplicity of presentation, we first consider the case of scalar measurements. In this setting, the optimal causal linear estimator of the state sequence $x_{t}$, in terms of mean-square error, is the well-known Kalman filter [11]. The procedure begins by specifying the initial state estimate and its corresponding error covariance matrix, denoted $\hat{x}_{0}^{+}$ and $\hat{P}_{0}^{+}$, respectively. (In our experiments, we initialize these as $\hat{x}_{0}^{+}=0$ and $\hat{P}_{0}^{+}=I_{n\times n}$.) Subsequent estimates and covariances are computed recursively using the Kalman filter’s prediction and update steps, as detailed by the expressions below.
@@
To examine whether a transformer can in-context learn to approximate the behavior of the Kalman filter, we design a training procedure based on synthetic data generated from linear dynamical systems with scalar observations. For each training instance, we randomly sample system parameters – including the state transmition matrix $F$, the measurement vectors $h_{1},\dots,h_{N}$, the process noise covariance $Q$, and the measurement noise covariance $\sigma^{2}$ – along with the corresponding outputs $y_{1},\dots,y_{N-1}$. These quantities are arranged into a structured input matrix of dimensions $(n+1)\times(2n+2N+1)$, formatted as
@@
As shown by Akyürek et al. [1], there exists a parametrization of a transformer’s head that can approximate the Read–Arithmetic–Write (RAW) operator defined in equations (11)–(12). Building on this, we identify a set of basic matrix operations, each implementable using the RAW operator, that are sufficient to reformulate the Kalman filter’s prediction and update steps. These operations are defined over specific index subsets of the transformer’s input matrix. For example, consider the system matrix $F$; the indices corresponding to $F$ in the input matrix of expression (51) are denoted by
@@
These buffers provide writeable memory slots for storing the evolving Kalman variables (such as the state, covariance matrix, and intermediate expressions) through the course of the algorithm. The complete recursive implementation is detailed in Algorithm 1. A line-by-line explanation of how each transformer operation maps to a Kalman step is available in Appendix A.
The above framework naturally extends to systems with vector-valued observations and uncorrelated noise. Suppose $y_{t}\in R^{m}$ and $r_{t}\sim\mathcal{N}(0,R)$, where $R$ is a diagonal matrix with positive entries $\sigma_{1}^{2}$, $\sigma_{2}^{2}$, $\dots$, $\sigma_{m}^{2}$. Let $H_{t}$ denote the measurement matrix, $y_{t}^{j}$ denote the $j^{th}$ component of $y_{t}$, and $H_{t}^{(j)}$ denote the $j^{th}$ row of $H_{t}$. Then the Kalman update equations become [10]
$\hat{x}_{t}^{(1)+}$ $=\hat{x}_{t}^{-}+\frac{\hat{P}_{t}^{-}H_{t}^{(1)T}}{H_{t}^{(1)}\hat{P}_{t}^{-}H_{t}^{(1)T}+\sigma_{1}^{2}}(y_{t}^{(1)}-H_{t}^{(1)T}\hat{x}_{t}^{-})$ (22)
$\hat{P}_{t}^{(1)+}$ $=(I-\frac{\hat{P}_{t}^{-}H_{t}^{(1)T}H_{t}^{(1)}}{H_{t}^{(1)}\hat{P}_{t}^{-}H_{t}^{(1)T}+\sigma_{1}^{2}})\hat{P}_{t}^{-}$ (23)
$\hat{x}_{t}^{(j)+}$ $=\hat{x}_{t}^{(j-1)+}+$
$\frac{\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}}{H_{t}^{(j)}\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}+\sigma_{j}^{2}}(y_{t}^{(j)}-H_{t}^{(j)T}\hat{x}_{t}^{(j-1)+})$
$j=2,\dots,m$ (24)
$\hat{P}_{t}^{(j)+}$ $=(I-\frac{\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}H_{t}^{(j)}}{H_{t}^{(j)}\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}+\sigma_{j}^{2}})\hat{P}_{t}^{(j-1)+}$
$j=2,\dots,m$ (25)
@@
This recursive structure enables sequential updates for each measurement dimension and is readily encoded in a format similar to that in the scalar case. In particular, the in-context learning can be performed by providing to the transformer the input formatted as
@@
Initialize $I_{\hat{X}_{\text{Curr}}}\leftarrow(1:n,\ 2n)$ ;
for $i=1$ to NN do
 $I_{\hat{X}_{\text{next}}}\leftarrow(1:n,\ 2n+2i)$ ;
 $I_{h}\leftarrow(1:n,\ 2n+2i-1)$ ;
 $I_{y}\leftarrow(0,\ 2n+2i)$ ;

### v3

@@ 1 Introduction
In-context learning (ICL), particularly in the form of few-shot prompting (Yogatama et al., 2019), has emerged as a prominent research direction in natural language processing (NLP). In this framework, a large language model (LLM) performs new tasks by conditioning on a small number of input–output examples provided in the prompt. One of the earliest works to demonstrate this capability was by Brown et al. (2020), who evaluated GPT-3 across a range of NLP datasets under “zero-shot", “one-shot" and “few-shot" settings. Zhao et al. (2021) identified majority-label bias, recency bias, and common-token bias as sources of instability in GPT-3’s accuracy under few-shot prompting, and proposed contextual calibration as a remedy. A theoretical explanation for ICL as a form of implicit Bayesian inference is offered in Xie et al. (2021). Min et al. (2022) explore why ICL works in practice, showing that its success is not dependent on having access to ground-truth labels; it further emphasized the importance of label space, input distribution, and prompt format. In parallel, Schlag et al. (2021) provide a theoretical analysis showing that transformers can act as fast weight programmers.
@@
Early work on using standard transformer decoders for in-context learning of autoregressive models includes (Garg et al., 2022), where the ability of transformers to learn function classes from examples is empirically investigated. A model is said to learn a function class $\mathcal{F}$ over a domain $\mathcal{X}$ if, for any $f\in\mathcal{F}$ and any inputs $x_{1},x_{2},\ldots,x_{N},x_{\text{query}}$ drawn i.i.d. from $\mathcal{X}$, the model can predict $f(x_{\text{query}})$ given the sequence $x_{1},f(x_{1}),x_{2},f(x_{2}),\ldots,x_{N},f(x_{N}),x_{\text{query}}$. The function classes studied by Garg et al. (2022) range from simple linear functions to sparse linear models, two-layer neural networks, and decision trees.
@@
Two parallel works by Von Oswald et al. (2023) and Akyürek et al. (2023) examine which algorithms transformers implicitly implement when learning function classes in-context. Von Oswald et al. (2023) build on Schlag et al. (2021) and show that linear self-attention layers can be interpreted as performing a step of gradient descent. Specifically, they prove that for a single-headed linear attention layer, there exist key, query, and value matrices such that a forward pass corresponds to one step of gradient descent with an $\ell_{2}$ loss applied to each token. In contrast, Akyürek et al. (2023) introduce a raw operator framework that supports operations such as matrix multiplication, scalar division, and memory-like read/write. They demonstrate that a single transformer head equipped with suitable key, query, and value matrices can approximate this operator, implying that transformers are, in principle, capable of performing linear regression via either stochastic gradient descent or in closed-form.
@@
Our work investigates whether transformers, using in-context learning (ICL), can learn to perform filtering in dynamical systems described by noisy state-space models. Specifically, we study whether a transformer, conditioned on a short context of past input–output pairs (and, optionally, system parameters), can predict the current output of the system. The model is pre-trained on synthetic trajectories generated by randomly sampled system parameters and evaluated without gradient-based updates at test time. Our goal is to understand which classical filtering algorithms the transformer’s behavior most closely resembles under this setup. To that end, we begin by examining the structure of the Kalman filter (Kalman, 1960), the optimal linear estimator for systems with Gaussian noise, and ask whether its operations can be replicated by a transformer architecture. We show that Kalman filtering can indeed be expressed using operations that are readily implemented by a transformer, and provide both a proof-by-construction and empirical evidence that, when sufficiently scaled and provided with enough context, the transformer learns to emulate Kalman-like behavior. Notably, its performance remains strong even when some parameters, such as the state-transition matrix, are omitted from the context, suggesting robustness and an ability to infer missing information implicitly. We then turn our attention to nonlinear systems and empirically demonstrate that transformers can in-context learn to achieve output prediction accuracy comparable to that of classical nonlinear filters such as the Extended Kalman Filter and particle filters. While our analysis and experiments primarily focus on output prediction, we also include preliminary evidence that transformers can, in some cases, explicitly recover latent states, thus highlighting their potential for more general inference tasks.
Scope and framing.
This paper focuses on the empirical investigation of whether transformers trained via in-context learning can perform inference in dynamical systems. In particular, our goal is to examine whether standard transformers, when trained on synthetic input–output trajectories from randomly generated systems, can learn to perform filtering tasks without explicit model supervision or architectural modifications. The study is guided by constructive arguments about the representational capacity of transformers. While we do not offer convergence guarantees, our results show that with sufficient size and context, transformers can approximate Kalman-like behavior in linear settings and extend to nonlinear regimes, sometimes outperforming classical filters. Performance may degrade under limited model capacity or severely reduced context, as discussed in Section 4.
@@ 1.1 Related work
Prior research at the interface of deep learning and Kalman filtering include Deep Kalman Filters (Krishnan et al., 2015) and KalmanNet (Revach et al., 2021), which learns the Kalman gain in a data-driven manner using a recurrent neural network. A follow-up study (Revach et al., 2022) extends this approach using gated recurrent units to estimate both the Kalman gain and noise statistics while assuming fixed system parameters. In contrast, our model is trained on data generated from randomly sampled system parameters, encouraging the transformer to learn the filtering procedure itself rather than memorize input–output mappings for a specific system. Other works have explored connections between attention mechanisms and structured state-space models (SSMs). Dao and Gu (2024) reformulate SSM computations as matrix multiplications on structured matrices, drawing parallels with efficient attention variants. Sieber et al. (2024) propose a unified dynamical systems framework encompassing attention, SSMs, RNNs, and LSTMs. However, neither of these works address state estimation, filtering, or in-context learning. Goel and Bartlett (2024) show that softmax self-attention can approximate the Nadaraya–Watson kernel smoother, which in turn resembles the Kalman filter. In contrast, our work directly targets in-context learning for dynamical systems and builds on the raw operator framework of Akyürek et al. (2023) to show that transformers can implement the exact operations needed for Kalman filtering, as supported by both theoretical constructions and empirical validation.
@@ 2.1 Transformers
Transformers, introduced by Vaswani et al. (2017), are a class of neural network architectures that rely on an attention mechanism to capture relationships among elements in an input sequence. This mechanism enables transformers to model long-range dependencies and has been central to their success in sequence-to-sequence tasks. The experiments in this paper utilize the GPT-2 architecture (Radford et al., 2019), a decoder-only variant of the transformer.
@@
where $\lambda$ denotes layer normalization and $\sigma$ is the activation function. In our experiments, we use the Gaussian Error Linear Unit (GeLU) activation (Hendrycks and Gimpel, 2016).
@@ 2.2 State-space models for linear dynamical systems
State-space models are widely used across machine learning (Gu et al., 2021), computational neuroscience (Barbieri et al., 2004), control theory (Kailath, 1980), signal processing (Kailath et al., 2000), and economics (Zeng and Wu, 2013). A central problem in many of these domains is the estimation of the latent state sequence $x_{t}$} and its functions (e.g., the system’s output) given noisy observations $y_{t}$} and the parameters of the model.
@@ 2.3 In-context learning in the absence of dynamics
  • Stochastic Gradient Descent. Initialize with $\hat{x}_{0}=0_{n\times 1}$, and update recursively as
    $\hat{x}_{t}=\hat{x}_{t-1}-2\alpha(h_{t-1}^{T}h_{t-1}\hat{x}_{t-1}-h_{t-1}^{T}y_{t-1}),$ (8)
@@
  • Ordinary Least Squares (OLS). Form the matrix $\bar{H}\in\mathbb{R}^{N\times n}$ with rows $h_{1},h_{2},\dots,h_{N}$, and let $\bar{Y}=[y_{1},y_{2},...,y_{N}]^{T}$. The OLS estimate is
    $\hat{x}_{OLS}=(\bar{H}^{T}\bar{H})^{-1}\bar{H}^{T}\bar{Y}.$ (9)
@@
    $\hat{x}_{Ridge}=(\bar{H}^{T}\bar{H}+\lambda I_{n\times n})^{-1}\bar{H}^{T}\bar{Y},$ (10)
@@
A pioneering study that explored the ability of language models to learn linear functions and implement simple algorithms was presented by Garg et al. (2022). Building on that work, Akyürek et al. (2023) examined whether GPT-2–based transformers can in-context learn the setting where $x_{t}=x$ as in (7). Specifically, they investigated which algorithms the transformer implicitly learns to implement when tasked with predicting $y_{N}$, given an input formatted as the matrix
@@
In Akyürek et al. (2023), the transformer was trained on batches of examples constructed from randomly sampled latent states and measurement parameters. The authors observed that the model’s behavior depends on both the architecture size and the context length. For small models trained with short contexts, the transformer approximates stochastic gradient descent (SGD). For moderate context lengths (up to the dimensionality of the latent state) and moderately sized models, its behavior resembles ridge regression. Finally, with large architectures and context lengths exceeding the latent dimension, the transformer’s performance approaches that of ordinary least squares (OLS).
A major contribution of Akyürek et al. (2023) was to theoretically demonstrate that transformers can approximate the operations required to implement SGD and closed-form regression. This was accomplished with the RAW (Read–Arithmetic–Write) operator (see Appendix G), parameterized by matrices $W_{o}$, $W_{a}$, and $W$, and with an element-wise operator $\circ\in\{+,*\}$. The RAW operator maps the input to layer $l$, denoted $G^{(l)}$, to the output $G^{(l+1)}$ using index sets $s$, $r$, $w$, a time map $K$, and token positions $i=1,\dots,2N$ as
$G_{i,w}^{(l+1)}$ $=W_{o}\left(\left[\frac{W_{a}}{|K(i)|}\sum_{k\in K(i)}G_{k}^{(l)}[r]\right]\circ WG_{i}^{(l)}[s]\right),$ (11)
$G_{i,j\notin w}^{(l+1)}$ $=G_{i,j\notin w}^{(l)}.$ (12)
A single transformer head can approximate this operator for arbitrary $W_{o}$, $_{a}, WW$, and $\circ$. With specific parameterizations, it can implement primitives such as affine transformations, matrix multiplication, scalar division, dot products, and memory-based read/write operations.
@@ 3 In-Context Learning for Dynamical Systems
We begin by outlining an in-context learning procedure for the generic linear state-space model defined in (4)–(5), assuming a time-invariant state equation (i.e., $F_{t}=F\neq I$, $Q\neq 0$). For simplicity of presentation, we first consider the case of scalar measurements. In this setting, the optimal causal linear estimator of the state sequence $x_{t}$, in terms of mean-square error, is the well-known Kalman filter (Kalman, 1960). The procedure begins by specifying the initial state estimate and its corresponding error covariance matrix, denoted $\hat{x}_{0}^{+}$ and $\hat{P}_{0}^{+}$, respectively. (In our experiments, we initialize these as $\hat{x}_{0}^{+}=0$ and $\hat{P}_{0}^{+}=I_{n\times n}$.) Subsequent estimates and covariances are computed recursively using the Kalman filter’s prediction and update steps, as detailed by the expressions below.
@@
To examine whether a transformer can in-context learn to approximate the behavior of the Kalman filter, we design a training procedure based on synthetic data generated from linear dynamical systems with scalar observations. For each training instance, we randomly sample system parameters – including the state transition matrix $F$, the measurement vectors $h_{1},\dots,h_{N}$, the process noise covariance $Q$, and the measurement noise covariance $\sigma^{2}$ – along with the corresponding outputs $y_{1},\dots,y_{N-1}$. These quantities are arranged into a structured input matrix of dimensions $(n+1)\times(2n+2N+1)$, formatted as
@@
As shown by Akyürek et al. (2023), there exists a parametrization of a transformer’s head that can approximate the Read–Arithmetic–Write (RAW) operator defined in equations (11)–(12). Building on this, we identify a set of basic matrix operations, each implementable using the RAW operator, that are sufficient to reformulate the Kalman filter’s prediction and update steps. These operations are defined over specific index subsets of the transformer’s input matrix. For example, consider the system matrix $F$; the indices corresponding to $F$ in the input matrix of expression (20) are denoted by
@@
These buffers provide writeable memory slots for storing the evolving Kalman variables (such as the state, covariance matrix, and intermediate expressions) through the course of the algorithm. The complete recursive implementation is detailed in Algorithm 1. A line-by-line explanation of how each transformer operation maps to a Kalman step is available in Appendix A. Note that Algorithm 1 serves as a symbolic abstraction describing how a generic transformer architecture can, in principle, emulate a Kalman filter using the Mul, Div, Aff operators that are special instances of the unified transformer primitive RAW operator. These derivations are not convergence or learning guarantees, but rather proofs of representability illustrating what the transformer architecture can compute in principle. The steps outlined in Appendix A illuminate this derivation in detail and walk through the mathematics behind each symbolic update. The empirical investigation presented later in this paper uses standard transformers, without any custom architectural modifications.
The above framework naturally extends to systems with vector-valued observations and uncorrelated noise. Suppose $y_{t}\in R^{m}$ and $r_{t}\sim\mathcal{N}(0,R)$, where $R$ is a diagonal matrix with positive entries $\sigma_{1}^{2}$, $\sigma_{2}^{2}$, $\dots$, $\sigma_{m}^{2}$. Let $H_{t}$ denote the measurement matrix, $y_{t}^{j}$ denote the $j^{th}$ component of $y_{t}$, and $H_{t}^{(j)}$ denote the $j^{th}$ row of $H_{t}$. Then the Kalman update equations become (Kailath et al., 2000)
$\hat{x}_{t}^{(1)+}$ $=\hat{x}_{t}^{-}+\frac{\hat{P}_{t}^{-}H_{t}^{(1)T}}{H_{t}^{(1)}\hat{P}_{t}^{-}H_{t}^{(1)T}+\sigma_{1}^{2}}\bigg(y_{t}^{(1)}-H_{t}^{(1)T}\hat{x}_{t}^{-}\bigg)$ (22)
$\hat{P}_{t}^{(1)+}$ $=\bigg(I-\frac{\hat{P}_{t}^{-}H_{t}^{(1)T}H_{t}^{(1)}}{H_{t}^{(1)}\hat{P}_{t}^{-}H_{t}^{(1)T}+\sigma_{1}^{2}}\bigg)\hat{P}_{t}^{-}$ (23)
$\hat{x}_{t}^{(j)+}$ $=\hat{x}_{t}^{(j-1)+}+\frac{\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}}{H_{t}^{(j)}\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}+\sigma_{j}^{2}}\bigg(y_{t}^{(j)}-H_{t}^{(j)T}\hat{x}_{t}^{(j-1)+}\bigg)\ \ \ \ j=2,\dots,m$ (24)
$\hat{P}_{t}^{(j)+}$ $=\bigg(I-\frac{\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}H_{t}^{(j)}}{H_{t}^{(j)}\hat{P}_{t}^{(j-1)+}H_{t}^{(j)T}+\sigma_{j}^{2}}\bigg)\hat{P}_{t}^{(j-1)+}\ \ \ \ j=2,\dots,m$ (25)
@@
While equations (13)-(17) present the standard Kalman filter in its classical recursive form, equations (22)-(27) re-express the same Kalman filter updates in an unrolled format that makes them amenable to transformer implementation. This recursive structure enables sequential updates for each measurement dimension and is readily encoded in a format similar to that in the scalar case. In particular, the in-context learning can be performed by providing to the transformer the input formatted as
@@
1 Initialize $I_{\hat{X}_{\text{Curr}}}\leftarrow(1:n,\ 2n)$ ;
2
3for $i=1$ to NN do
4 $I_{\hat{X}_{\text{next}}}\leftarrow(1:n,\ 2n+2i)$ ;
5 $I_{h}\leftarrow(1:n,\ 2n+2i-1)$ ;
6 $I_{y}\leftarrow(0,\ 2n+2i)$ ;
7